# Conformal Off-Policy Prediction for Multi-Agent Systems

### v1

@@ Abstract
Off-Policy Prediction (OPP), i.e., predicting the outcomes of a target policy using only data collected under a nominal (behavioural) policy, is a paramount problem in data-driven analysis of safety-critical systems where the deployment of a new policy may be unsafe. To achieve dependable off-policy predictions, recent work on Conformal Off-Policy Prediction (COPP) leverage the conformal prediction framework to derive prediction regions with probabilistic guarantees under the target process. Existing COPP methods can account for the distribution shifts induced by policy switching, but are limited to single-agent systems and scalar outcomes (e.g., rewards). In this work, we introduce MA-COPP, the first conformal prediction method to solve OPP problems involving multi-agent systems, deriving joint prediction regions for all agents’ trajectories when one or more “ego” agents change their policies. Unlike the single-agent scenario, this setting introduces higher complexity as the distribution shifts affect predictions for all agents, not just the ego agents, and the prediction task involves full multi-dimensional trajectories, not just reward values. A key contribution of MA-COPP is to avoid enumeration or exhaustive search of the output space of agent trajectories, which is instead required by existing COPP methods to construct the prediction region. We achieve this by showing that an over-approximation of the true JPR can be constructed, without enumeration, from the maximum density ratio of the JPR trajectories. We evaluate the effectiveness of MA-COPP in multi-agent systems from the PettingZoo library and the F1TENTH autonomous racing environment, achieving nominal coverage in higher dimensions and various shift settings.
@@ I Introduction
To obtain reliable predictions, we leverage the framework of conformal prediction (CP) [5, 6], a technique to derive prediction regions guaranteed to cover the true (unknown) output with arbitrary probability. Crucially, these probabilistic guarantees are finite-sample (non-asymptotic) and distribution-free, i.e., they don’t rely on priors or parametric assumptions about the data distribution. For these properties, it is unsurprising that CP has become, in recent years, the go-to method for uncertainty quantification, especially in safety-critical applications, see, e.g., [7, 8, 9, 10].
CP uses a set of calibration points to derive a distribution of model residuals, aka scores. The prediction region for a test point is then obtained by including all outputs whose scores appear sufficiently likely w.r.t. such calibration distribution. This procedure yields the desired probabilistic guarantees as long as the calibration and test data are exchangeable (a weaker assumption than i.i.d.). Exchangeability, however, is violated in the presence of distribution shifts, which are inherent to OPP. The framework of weighted exchangeability [11] extends CP to handle distribution shifts by reweighting the calibration points “as if” they were sampled under the target distribution. This is achieved by using estimates of the density ratios (DRs) between target and behavioural distributions. Recent work on Conformal Off-Policy Prediction (COPP) [12, 13, 14] leverage CP and weighted exchangeability to provide valid prediction regions for OPP problems. These methods, however, only support single-agent systems and construct regions for scalar outcomes (e.g., reward).
@@
Indeed, constructing a JPR valid under generic distribution shifts using CP and weighted exchangeability normally requires, for each test input, an exhaustive search over the output trajectory space, which is infeasible in our high-dimensional settings. To overcome this problem, we demonstrate that we can derive a conservative over-approximation of the true (unknown) JPR without resorting to exhaustive search, if the maximum density ratio (DR) over all the JPR’s trajectories is known. Building on this result, in MA-COPP, we pivot the search task over the maximum DR, which is significantly more effective than an exhaustive output search.
@@
Figure 1: 2D visualisation of JPRs constructed over the position of a single ego agent
@@ II The Multi-Agent OPP Problem
We can observe a set of $N$ state trajectories of length $T$, and for ego agents only, we can also observe the actions they perform at each step of the trajectory. More formally, we have access to the following dataset of realisations of the process $\left(\mathbf{X}_{1},\left(A_{e,1}\right)_{e\in E},\mathbf{X}_{2},\left(A_{e,2}\right)_{e\in E},\ldots,\mathbf{X}_{T}\right)$ induced by (1):
@@ III Background
    Note that the above holds in the more general case when $(x,y)$ is exchangeable w.r.t. calibration data.
@@ IV The MA-COPP approach
For simplicity, we denote $\mathbf{X}_{1\ldots H}$ by $X$, $\mathbf{X}_{H+1\ldots T}$ by $X^{\prime}$, and $\mathbf{A}_{H+1\ldots T,e}$ by $A$. So, we can define the coverage of $(X,X^{\prime},A)$ (the augmented sequence) viz. the prediction region $C_{\alpha}^{+}(X)$ by $\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]$, where $f$ is the indicator function telling if $(X^{\prime},A)\in C_{\alpha}^{+}(X)$. By the premises of the proposition, we have $\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]\geq 1-\alpha$. So, we need to prove that $\mathbb{E}_{X,X^{\prime}}[g(X,X^{\prime})]\geq\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]$ where $g$ is the indicator function telling if $X^{\prime}\in\Pi_{x^{\prime}}(C_{\alpha}^{+}(X))$. We have the following derivation:
@@ VI Related Work
Past COPP work and our work leverage the weighted exchangeability method of [11] (discussed in more detail in Section III), but other CP-based approaches exist that address the distribution shift problem, such as adaptive CP [31], an approach that dynamically adjusts the coverage level in an online manner to compensate for observed under/over-coverage under unknown distribution shifts; or, robust CP [32], which extends CP to handle bounded adversarial input perturbations.

### v2

@@ Abstract
Off-Policy Prediction (OPP), i.e., predicting the outcomes of a target policy using only data collected under a nominal (behavioural) policy, is a paramount problem in data-driven analysis of safety-critical systems where the deployment of a new policy may be unsafe. To achieve dependable off-policy predictions, recent work on Conformal Off-Policy Prediction (COPP) leverage the conformal prediction framework to derive prediction regions with probabilistic guarantees under the target process. Existing COPP methods can account for the distribution shifts induced by policy switching, but are limited to single-agent systems and scalar outcomes (e.g., rewards). In this work, we introduce MA-COPP, the first conformal prediction method to solve OPP problems involving multi-agent systems, deriving joint prediction regions for all agents’ trajectories when one or more “ego” agents change their policies. Unlike the single-agent scenario, this setting introduces higher complexity as the distribution shifts affect predictions for all agents, not just the ego agents, and the prediction task involves full multi-dimensional trajectories, not just reward values. A key contribution of MA-COPP is to avoid enumeration or exhaustive search of the output space of agent trajectories, which is instead required by existing COPP methods to construct the prediction region. We achieve this by showing that an over-approximation of the true joint prediction region (JPR) can be constructed, without enumeration, from the maximum density ratio of the JPR trajectories. We evaluate the effectiveness of MA-COPP in multi-agent systems from the PettingZoo library and the F1TENTH autonomous racing environment, achieving nominal coverage in higher dimensions and various shift settings.
@@ I Introduction
To obtain reliable predictions, we leverage the framework of conformal prediction (CP) [5, 6], a technique to derive prediction regions guaranteed to cover the true (unknown) output with arbitrary probability. Crucially, these probabilistic guarantees are finite-sample (non-asymptotic) and distribution-free, i.e., they don’t rely on priors or parametric assumptions about the data distribution. For these properties, it is unsurprising that CP has become, in recent years, the go-to method for uncertainty quantification, especially in safety-critical applications, see, [7, 8, 9, 10].
CP uses a set of calibration points to derive a distribution of model residuals, a.k.a. scores. The prediction region for a test point is then obtained by including all outputs whose scores appear sufficiently likely w.r.t. such calibration distribution. This procedure yields the desired probabilistic guarantees as long as the calibration and test data are exchangeable (a weaker assumption than i.i.d.). Exchangeability, however, is violated in the presence of distribution shifts, which are inherent to OPP. The framework of weighted exchangeability [11] extends CP to handle distribution shifts by reweighting the calibration points “as if” they were sampled under the target distribution. This is achieved by using estimates of the density ratios (DRs) between target and behavioural distributions. Recent works on Conformal Off-Policy Prediction (COPP) [12, 13, 14] leverage CP and weighted exchangeability to provide valid prediction regions for OPP problems. These methods, however, only support single-agent systems and construct regions for scalar outcomes (e.g., reward).
@@
Indeed, constructing a valid JPR under generic distribution shifts using CP and weighted exchangeability normally requires, for each test input, an exhaustive search over the output trajectory space, which is infeasible in our high-dimensional settings. To overcome this problem, we demonstrate that we can derive a conservative over-approximation of the true (unknown) JPR without resorting to exhaustive search, if the maximum density ratio (DR) over all the JPR’s trajectories is known. Building on this result, in MA-COPP, we pivot the search task over the maximum DR, which is significantly more effective than an exhaustive output search.
@@
Figure 1: 2D visualisation of actual JPRs constructed over the position of a single ego agent in the MPE environment defined in § V-B1
@@ II The Multi-Agent OPP Problem
We can observe a set of $N$ global state trajectories of length $T$, and for ego agents only, we can also observe the actions they perform at each step of the trajectory. More formally, we have access to the following dataset of realisations of the process $\left(\mathbf{X}_{1},\left(A_{e,1}\right)_{e\in E},\mathbf{X}_{2},\left(A_{e,2}\right)_{e\in E},\ldots,\mathbf{X}_{T}\right)$ induced by (1):
@@ III Background
    Note that the above holds in the more general case when $(x,y)$ is exchangeable w.r.t. calibration data, i.e., when the joint probability of $(x^{1},y^{1}),\ldots,(x^{|\mathcal{D}_{c}|},y^{|\mathcal{D}_{c}|}),(x,y)$ remains the same for any permutation of the data points.
@@ IV The MA-COPP approach
For simplicity, we denote $\mathbf{X}_{1\ldots H}$ by $X$, $\mathbf{X}_{H+1\ldots T}$ by $X^{\prime}$, and $\mathbf{A}_{H+1\ldots T,e}$ by $A$. So, we can define the coverage of $(X,X^{\prime},A)$ (the augmented sequence) w.r.t. the prediction region $C_{\alpha}^{+}(X)$ by $\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]$, where $f$ is the indicator function telling if $(X^{\prime},A)\in C_{\alpha}^{+}(X)$. By the premises of the proposition, we have $\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]\geq 1-\alpha$. So, we need to prove that $\mathbb{E}_{X,X^{\prime}}[g(X,X^{\prime})]\geq\mathbb{E}_{X,X^{\prime},A}[f(X,X^{\prime},A)]$ where $g$ is the indicator function telling if $X^{\prime}\in\Pi_{x^{\prime}}(C_{\alpha}^{+}(X))$. We have the following derivation:
@@ VI Related Work
Past COPP work and our work leverage the weighted exchangeability method of [11] (discussed in more detail in Section III), but other CP-based approaches exist that address the distribution shift problem, such as adaptive CP [31], an approach that dynamically adjusts the coverage level in an online manner to compensate for observed under/over-coverage under unknown distribution shifts, or robust CP [32], which extends CP to handle bounded adversarial input perturbations.
@@ VII Conclusion
VIII Disclosure of Funding
This work is supported by the REXASI-PRO H-EU project, call HORIZON-CL4-2021-HUMAN- 01-01, grant agreement ID: 101070028; the Engineering and Physical Sciences Research Council (EPSRC) under Award EP/W014785/2; NSF CCRI 1925587; NSF GRFP award DGE-2236662; and US DoT Safety21 National University Transportation Center.